### v1

# JEPA-Reasoner: Decoupling Latent Reasoning from Token Generation

Abstract
While Joint-Embedding Predictive Architecture (JEPA) has emerged as a powerful architecture for learning rich latent representations, it fundamentally lacks generative abilities. Meanwhile, latent space reasoning attempts for Transformer models like COCONUT do improve performance, but they ultimately rely on token-by-token generation, which still accumulates compounding error and relies on context information to gain reasoning insights. To address these limitations, we propose JEPA-Reasoner, a novel JEPA model enhanced with generative ability that reasons in latent space. We augment it with a separate action-taker model, Talker, to produce human-readable sentences. Our approach demonstrates that decoupling latent space reasoning and token generation enables JEPA-Reasoner to produce mixed latent vectors that might lay the foundation for multi-threaded reasoning, while performing autoregressive generation with superior robustness to compounding error.
Foundation Models, Latent Reasoning, ICML
1 Introduction
The Joint-Embedding Predictive Architecture (JEPA) (i-jepa) has demonstrated strong performance in learning semantic world representations, exhibiting superior world understanding ability compared with traditional end-to-end generative models that work in pixel or token space. By predicting abstract representations in latent space, the JEPA architecture is able to filter irrelevant details and preserve essential information needed for prediction (i-jepa). Such architecture has been proven to be a viable approach for representation learning and foundation model development. Various JEPA implementations, including I-JEPA (i-jepa), V-JEPA2 (v-jepa2), and M3-JEPA (m3-jepa), have shown success across various modalities and downstream tasks.
However, JEPA models are inherently non-generative (m3-jepa) because of their objective: filling missing information in the current state (i-jepa), rather than generating new content. Besides, the predictor often requires a detailed, predetermined target state or task instruction (e.g., V-JEPA 2-AC (v-jepa2)), which is often unavailable in tasks demanding long-term planning and step-by-step reasoning. This feature limits the application of the broad knowledge in JEPA models to generative tasks.
Furthermore, while traditional token-level autoregressive models have sequential reasoning capabilities, their token-by-token generation process is prone to compounding errors. Even if the predicted probabilities of tokens are weighted and combined, LLMs still cannot go beyond single-threaded reasoners (llms-reason-singel-thread). Although several previous research have explored latent space reasoning for Transformer models, the end-to-end generation goal of a single coupled model limited the full potential of latent space reasoning, while also making training complex and inefficient (coconut).
To address these limitations, we propose JEPA-Reasoner, a novel decoupled architecture that utilizes separate models for reasoning and token generation. The reasoning model, JEPA-Reasoner, transforms the JEPA framework from a target-conditioned system into an autoregressive generative model. Operating entirely within the continuous, normalized latent space, JEPA-Reasoner focuses solely on latent space reasoning, offloading the token generation task to its action-taker module: Talker. In contrast to existing latent space reasoning solutions like COCONUT, the division of objective frees the reasoner from expression burden and enables continuous latent guidance (compared with non-continuous latent guidance that relies on context information to retrieve latent reasoning results (coconut)) during the token generation process.
Our key insight is that performing pure reasoning in an abstract, continuous, and normalized latent space without token generation burden allows for the construction of high-level reasoning chains that carry rich semantic information throughout the autoregressive process while correctly ignoring irrelevant details or distracting information. Our empirical experiments show further improved abilities beyond that: by operating in the latent space, JEPA-Reasoner can maintain multiple hypotheses during the reasoning process simultaneously and mitigate the catastrophic error propagation associated with discrete token sampling.
2 Related Work
Joint-Embedding Predictive Architectures (JEPA).
JEPA (i-jepa) introduced a framework that makes predictions in representation space. It utilizes self-supervised training to learn latent states that are not directly human-readable. A predictor module was trained to predict the target state based on encoded inputs. Multiple variants of this architecture have extended the JEPA family to various modalities and downstream tasks (i-jepa). However, these models are non-generative. Attempts to make JEPA generative, such as D-JEPA (d-jepa), utilize the learned representations to condition diffusion models for data generation (e.g., text to images, text to audio) but still failed to enable sequential reasoning or planning within the JEPA framework itself. In contrast, our key innovation is to adapt the core JEPA objective for autoregressive latent-space generation.
Latent Space Reasoning.
Previous work on latent space reasoning mainly focuses on looping hidden states, either through horizontal autoregression like COCONUT (coconut), or vertical recurrent depth scaling (recurrent-latent-reasoning). However, these paradigms utilize a single coupled model for both latent reasoning and token generation, ignoring the mismatch between the two tasks: one requires high-level global planning, decision-making, and choice tracing, while the other requires local correctness in grammar and fluency. Besides, whether iterating over tokens or deepening the computation per token, both paradigms remain bound by causal masking, meaning early token generation cannot be guided by future reasoning states. Considering these limitations, our key innovation is to decouple latent space reasoning and token generation. By generating a complete latent chain first, we enable consistent latent guidance where all tokens are generated from the full reasoning trajectory. This produces an answer with higher quality that is less prone to error propagation. Additionally, the decoupled design enables efficient optimization with a single forward pass in latent space, unlike coupled models like COCONUT or Recurrent Depth transformers, which require multiple synchronized passes or complex recurrent unrolling.
Autoregressive Models and Robustness.
Modern Transformer models conduct token-level prediction in an autoregressive manner. While proven powerful on various tasks, this approach is known to suffer from compounding errors in long-horizon tasks. Techniques like Chain-of-Thought (cot) improve reasoning by generating intermediate steps, but still operate at the token level. JEPA-Reasoner aims to improve robustness of autoregressive generation by moving the reasoning process into a continuous, abstract latent space, reducing the impact of localized errors.
3 Model Architecture
JEPA-Reasoner decouples the reasoning process from output generation, making next-state predictions completely dependent on previously generated, semantic-rich, lossless latent states. The architecture consists of:
  • JEPA-Reasoner: Generate sequential latent space reasoning chains independently.
  • Talker: Translates the latent states into tokens. Note that the Talker is not able to make predictions. Its task is to reconstruct tokens based entirely on the latent output from JEPA-Reasoner.
3.1 JEPA-Reasoner
Figure 1: Architecture of JEPA-Reasoner and its action taker (Talker). The Reasoner consists of an embedding layer as the token encoder and Transformer blocks as the predictor. The embedding layer for input tokens always uses the latest weights, while the weight of the embedding layer for target tokens is the exponential moving average of the historical weights of the input embedding layer.
Model Components.
JEPA-Reasoner follows the JEPA philosophy, containing an embedding layer as a textual token encoder and modified Transformer blocks for the predictor, since the Transformer has proven its strong ability in sequence modeling. After the modified Transformer blocks, we applied a hybrid normalization layer (RMS and L2 normalization). We utilize L2 normalization to prevent exploding magnitude caused by residual connections. In the modified Transformer block, we apply a non-learnable QK-Norm (qk-norm) to make it more numerically stable.
Latent Space Generation.
Unlike traditional JEPA models, in which the predictor is aimed at filling missing information in the current state (i-jepa), the predictor of JEPA-Reasoner generates the next latent matrix representing the subsequent reasoning steps. Crucially, this generated latent matrix is not projected into vocabulary probabilities via an LM head. Instead, it is normalized by the hybrid normalization layer to the unit hypersphere and looped back as the input of the first Transformer block for the next round of autoregressive generation in the latent space.
Training Objective and Target Encoder.
The model is trained to predict latent representations provided by a target encoder. Following standard JEPA methodology, the target encoder weights are an exponential moving average (EMA) of the data encoder weights, providing stable and rich training targets. Given the normalized nature of our latent space, we use scaled cosine similarity loss computed entirely in latent space, ensuring the predictor learns consistent feature representations and dynamics (refer to Section 4.2 for more detail).
3.2 Action-Taker Model
In this experiment, there is only one action-taker model: Talker. The Talker model is a standard Transformer-based model trained independently. We designed two variants of Talker: Mono-Talker and Dual-Talker. Detailed information about the components of the two Talkers is shown in Table 1. Mono-Talker does not have an embedding layer or encoders, it only has decoders. Mono-Talker is designed for reconstruction tasks that do not require context information, receiving latent vectors from JEPA-Reasoner and constructs the complete token sequence in one forward pass. Dual-Talker is designed for context-aware reconstruction, usually necessary in natural language tasks. It has an embedding layer, encoders, and decoders. The embedding layer is used for encoding previously determined outputs of JEPA-Reasoner that contain contextual information, while the encoder blocks receive latent vectors from JEPA-Reasoner as input. The decoders generate tokens autoregressively conditioned on previous tokens, with continuous latent guidance from the output of encoders. However, Dual-Talker was trained for reconstruction rather than generation, as our ablation study (Appendix C) showed that Talker is critically dependent on the Reasoner’s output. During training, the JEPA-Reasoner is frozen. Talker receives the sequence of latent vectors and is trained to reconstruct the corresponding token sequence using standard cross-entropy loss.
Table 1: Components of Mono-Talker and Dual-Talker
Mono-Talker Dual-Talker
Embedding Layer No Yes
Standard Encoder No Yes
Standard Decoder Yes Yes
LM Head Yes Yes
4 Training Procedure
The training process of JEPA-Reasoner consists of two main phases. The first stage is pretraining that teaches basic knowledge (e.g., grammar and commonsense) to the model. The second stage is self-supervised training (SST), which adapts the model to perform consistent latent space reasoning.
4.1 Pretraining
We apply established Transformer training methods to provide the model with basic knowledge and language understanding capabilities.
Objective and Methodology.
The model is trained as a standard decoder-only Transformer on the next-token prediction task in a teacher-forcing way. We employ tied word embeddings that share the weight of the embedding layer with a temporary LM head. The LM head is only used in pretraining and is removed after pretraining is finished. The L2 normalization layer is disabled in the pretraining phase to enable simple and direct reuse of current Transformer training recipes. Considering that tied word embedding encourages $\mathbf{W}_{Embed}\cdot\mathbf{W}_{Embed}^{T}=I$ and $\mathbf{v}_{pred}\cdot\mathbf{v}_{embed}=\|\mathbf{v}_{pred}\|\cdot\|\mathbf{v}_{embed}\|\cdot\cos(\theta)$, the tied-embedding approach will indirectly encourage angular alignment between predicted vectors and embedding vectors, which facilitates the subsequent transition from token-level to latent-level prediction.
4.2 Self-Supervised Training
The SST phase adapts the pretrained model for making consistent predictions in the continuous latent space. Since the model is fully transforming into a latent space reasoner instead of a token generator, the ability to produce correct logits no longer matters. Considering this, we apply similar self-supervised training as Meta’s JEPA series (i-jepa) in this stage. Without the need for autoregressively generating final token outputs to compute a loss, self-supervised training enables efficient parallel training compared with COCONUT (coconut).
Objective and Methodology.
The temporary LM head in the pretraining stage was discarded and L2 normalization layer was restored. The model is now optimized to predict the latent representation of the next sequence segment, with a consistent dimensional semantics as what the embedding layer produces. We switch to scaled cosine distance loss, aligning with the L2 normalization strategy used to ensure stability during autoregressive looping and focusing the learning on angular differences:
$\mathcal{L}(\theta,\theta^{\prime})=k-k\cdot\text{cos}(h_{\text{pred}}(\theta),h_{\text{target}}(\theta^{\prime}))$ (1)
where $k$ is the scalar, $h_{\text{pred}}$ is the predicted latent vector by the Reasoner (parameters $\theta$), and $h_{\text{target}}$ is the target latent vector from the EMA encoder (parameters $\theta^{\prime}$). In our empirical tests, we find that normal cosine distance loss failed to support enough optimization when the loss is small. We tested a series of $k$ values, and chose $k=4$ in our experiments (Refer to Appendix D for more detail).
Target Generation.
The weight updating method of the target embedding layer is different from that of the input embedding layer. The input embedding layer always applies the latest weights, while the target embedding layer utilizes exponential moving average to generate its weights from the historical weights of the input embedding layer. We applied a high momentum value of 0.98 to prevent rank collapse in the embedding layer while ensuring enough space to adjust for angular alignment.
5 Latent Space Properties
We analyze the property of JEPA-Reasoner’s latent representation on two synthetic tasks designed to probe specific capabilities in controlled environments: mixed latent vector generation via a tree-search problem, and robustness to error propagation via a Context-Free Grammar (CFG) generation task.
5.1 Continuous Representation of Uncertainty
Within a reasoning process, JEPA-Reasoner is able to produce mixed latent vectors that are not limited to the discrete representations in the embedding layers. The mixed latent vectors approximate a linear combination of more than one vocabulary latent (latent vectors that correspond to individual vocabulary tokens). To systematically examine this behavior, we trained a smaller JEPA-Reasoner (42M) to search routes from the root to specific leaves in a binary tree.
5.1.1 Data Preparation
Training data consists of randomly generated binary trees with depth limited to 4. Each tree node was represented by a character with a unique token, making up the vocabulary along with other special tokens. In the generation process, we randomly pick node names to prevent the model from memorizing relationships based on names. Refer to Appendix A for an example.
5.1.2 Model Configuration
The JEPA-Reasoner model and Mono-Talker model were built with specifications stated in Table 2. We chose the combination of JEPA-Reasoner with Mono-Talker because this task does not require context-aware reconstruction.
Table 2: Model configurations in tree-search experiment
Latent Dim. Attention Dim. FFN Dim. Head Count Decoder Count
JEPA-Reasoner 384 768 1536 16 18
Mono-Talker 384 768 1536 8 6
5.1.3 Training
The pretraining and SST process are completely the same as stated in Section 4, except for loss masking. In the pretraining stage, loss was computed on all positions, while in SST, loss was only computed on the positions that define the desired route. When training the Talker model, we only passed latent vectors that describe the route to Talker to ensure it had no access to the tree structure or the target leaf, which guaranteed the Talker could not solve the task on its own.
5.1.4 Results and Conclusions
The final combination of the JEPA-Reasoner and Mono-Talker models achieved 99.87% accuracy (exact match) in searching routes from the tree root to specific leaves. Given the restricted context window of the Mono-Talker model, we could confirm that only JEPA-Reasoner was responsible for reasoning. Based on this result, we examine the generated latent vectors to probe the reasoning behavior of JEPA-Reasoner.
We calculated the distance from the predicted latent vector to the plane spanned by any two vocabulary vectors and sorted them from closest to farthest. In the sorted list, the plane spanned by latent vectors of sibling nodes frequently exhibits lower distances to the predicted latent vector, with an average ranking of top 1.72% in the ordered list. Also, we figured out all coefficient sets, $\alpha$ and $\beta$, that satisfy $\alpha\cdot\mathbf{l}_{0}+\beta\cdot\mathbf{l}_{1}=\mathbf{l}_{proj}$, where $\mathbf{l}_{0}$ and $\mathbf{l}_{0}$ are latent vectors of sibling nodes and $\mathbf{l}_{proj}$ is the projection of the predicted latent vector on the spanned plane. After comparing $\alpha$ and $\beta$, we find that for 99.89% of the time, the latent vector of the node on the correct route contributes more than the other sibling node. This discovery demonstrated that JEPA-Reasoner could make correct choices without completely discarding the other information that contains potentially correct choices. According to the previous COCONUT study (coconut), this behavior might lay the foundation for breadth-first multi-threaded reasoning.
5.2 Robustness to Error Perturbation
In the following sections, we demonstrate that decoupling reasoning chain generation from token production enables superior robustness under noisy conditions. While coupled models must simultaneously maintain reasoning coherence and produce correct tokens, our decoupled approach allows the reasoning model to focus solely on maintaining logical consistency in latent space, exhibiting better generation quality. This section contains two experiments that focus on two different sources of errors: token level error and latent space noise.
5.2.1 Experiment Methods
Robustness Test for Token Level Error
To evaluate the robustness of the decoupled model on token-level errors in the input sequence, we randomly replace 0% to 30% ground truth tokens in the input sequence with incorrect tokens. We compare the performance of JEPA-Reasoner and traditional Transformer models on multi-step completion tasks [^1] All scores obtained in these two tests are by testing the model across 5248 samples randomly chosen from the test dataset containing 100000 samples to minimize the bias introduced by randomness. using the exact match metric.
Robustness Test for Latent Space Error
To evaluate the robustness of the JEPA-Reasoner model on perturbations in latent space, we compare the performance of JEPA-Reasoner and the coupled continuous reasoning model COCONUT on multi-step completion tasks. In this experiment, we let the COCONUT model autoregressively generate 4 latent vectors first, followed by 4 tokens. For JEPA-Reasoner, we simply let it generate 8 latent vectors and use the Talker module to reconstruct 8 tokens. For both models, we add Gaussian noise to the generated latent vector at each step with $\mu=0$ and $\sigma$ ranging from 0% to 15% of the maximum value in the model’s output. Accuracy is calculated across the last 4 tokens with the exact match metric.
5.2.2 Data Preparation
Considering that the compounding error caused by different faulty tokens differs significantly, it is difficult to quantitatively analyze the model’s behavior under token-level errors (e.g., replacing keywords in the sentence will decrease the quality more considerably than replacing a word that functions as a connector). We followed previous work (allen-physics) and utilized Context-Free Grammar (CFG) for both experiments to create a controllable experiment setting.
Our custom CFG production rule features three terminal symbols with rule lengths of 3 or 4. With this rule, we generated long (approximately 600 to 700 symbols) and complex sequences that require non-trivial work to solve. The complexity of the grammar ensures that high accuracy relies on learning the underlying structure of the CFG sequence rather than memorizing specific sequences (refer to Appendix B.1 for full CFG specifications and production methods).
5.2.3 Model Configurations and Training Methods
We denote the vanilla Transformer baseline as $T$, the COCONUT-style coupled latent space reasoning model as $C$, and the decoupled model as $R$ (both JEPA-Reasoner and Talker are included). Table 3 shows detailed model configurations used in this section:
Table 3: Model Configurations for CFG Task. Talker Block Count format (E+D) refers to Encoder and Decoder blocks in the Dual-Talker model.
$R$ $T$ $C$
Total Parameters 315M 338M 338M
Latent Dimension 960 960 960
Attention Dimension 960 960 960
FFN Dimension 3840 3840 3840
Head Count 16 16 16
Talker Block Count 4 + 4
Reasoner Block Count 16
Transformer Block Count 24 24
Total Blocks 24 24 24
We first pretrain the Transformer models on CFG data using cross-entropy loss in token space. Since the pretraining methods of Transformer, COCONUT, and JEPA-Reasoner are identical, subsequent trainings are based on the same pretraining checkpoint [^2] Due to architectural differences, we initialize JEPA-Reasoner models using only the first $N$ blocks from the pretrained Transformer, where $N$ matches the JEPA-Reasoner’s block count.. We conduct post-training to obtain the Transformer baseline $T$. COCONUT models are trained to first predict 4 hidden states, then generate 4 tokens. Cross-entropy loss is computed between the output logits and the target sequence, excluding the hidden state positions. We follow the training method mentioned in Section 4 to train the JEPA-Reasoner and Dual Talker model $R$. For all training processes mentioned above, we apply identical hyperparameters (with learning rate of $1\times 10^{-4}$, effective batch size of 128, and context length of 1024) and train until their loss stabilizes, then checkpoints of best performance were chosen as the representative.
5.2.4 Results and Conclusions
Our robustness evaluation demonstrates the advantages of the decoupled architecture. In the token-level error experiment, JEPA-Reasoner demonstrated less performance degradation when facing input noise during multi-step CFG completion tasks (Figure 2). JEPA-Reasoner also exhibits higher performance across different magnitudes of Gaussian noise in the latent space error experiment (Table 4), providing more empirical evidence for its robustness advantage.
Figure 2: The relative performance of JEPA-Reasoner $R$ and Transformer baseline $T$ at each generation step with different fractions of correct tokens in the input sequence being replaced with wrong ones.
Table 4: Performance of JEPA-Reasoner $R$ and COCONUT $C$ under different latent noise levels.
$R$ $C$
$\sigma=0.0$ 0.4588 0.3740
$\sigma=0.05\times\max{(h_{t})}$ 0.4681 0.3688
$\sigma=0.10\times\max{(h_{t})}$ 0.4643 0.3650
$\sigma=0.15\times\max{(h_{t})}$ 0.4468 0.3629
These results demonstrate that JEPA-Reasoner has the potential to address the limitations of existing paradigms in Section 2. By operating in a normalized latent space and offloading token generation to the Talker module, subsequent reasoning outputs do not condition on previous decisions, thereby mitigating error accumulation in the autoregressive process and enabling more robust sequential generation under noisy conditions.
6 Real World Evaluation
We trained a 694M JEPA-Reasoner and a 198M Mono-Talker on natural language using the training steps stated in Section 4. We analyze the performance in two ways:
  1. 1.
    Performance difference between pretrained model (token level prediction) and JEPA-Reasoner (latent level prediction), paired with its Mono-Talker model.
  2. 2.
    Performance difference between JEPA-Reasoner (paired with Mono-Talker) and other latent or non-latent reasoning models.
6.1 Performance Gain After Latent Reasoning Adaptation
We evaluated both the base Transformer model and JEPA-Reasoner on GSM8K (gsm8k). Table 5 shows the benchmark score under both 5-shot and 8-shot settings.
Table 5: GSM8K benchmarks of base Transformer and JEPA-Reasoner
Model Accuracy (%)
5-shot 8-shot
Base Transformer 20.7 20.8
JEPA-Reasoner + Talker 39.4 51.9
Performance Gain.
Comparing to the base Transformer, JEPA-Reasoner improves performance by 79.2% (5-shot) and 149.5% (8-shot), respectively, demonstrating that adapting to a decoupled latent space reasoning paradigm significantly increases the accuracy of expressing internally learned knowledge with explicitly stated token sequences. Considering that it takes 300k steps for pretraining but only 13k steps for SST, while using the same dataset, it shows great evidence that most of the performance gain came from the adaptation to latent reasoning, rather than SST steps. Also, the scaled cosine similarity loss used in SST encourages smooth latent representation transition rather than logical correctness, which further support that minimal new knowledge was learned during SST.
In Context Learning.
Crucially, while the Base Transformer’s performance stagnates between 5-shot and 8-shot settings (improving only by 0.1%), the JEPA-Reasoner utilizes the additional examples to nearly double its accuracy. This scaling behavior indicates that the decoupled latent space reasoning architecture effectively overcomes the reasoning plateau, which was widely observed in small-scale token-based models, allowing for more robust logical deduction that is less constrained by surface-level token statistics.
6.2 Performance Comparison with Other Reasoning Models
We compare JEPA-Reasoner against standard Transformer models and other reasoning models. Table 6 shows the GSM8K benchmark performance of our JEPA-Reasoner model (paired with Talker) compared to other models of similar size or performance.
Table 6: GSM8K benchmark comparison between JEPA-Reason and other reasoning models. Note that we could not find 8-shot GSM8K benchmark results for COCONUT models. Since COCONUT models sacrifice performance (compared to CoT) as a trade-off for efficiency (coconut), we did not include them in the performance comparison.
Reasoning Paradigm Example Model Size 8-shot Accuracy (%)
Standard Gemma 3 4B 38.4
CoT Llama 3.2 1B 44.4
CoT Qwen 3 0.6B 42.5
Recurrent Depth Huginn-0125 3.5B 42.1
Ours JEPA-Reasoner 0.9B 51.9
Analysis of Results.
As illustrated in Table 6, JEPA-Reasoner demonstrates superior performance on the GSM8K benchmark compared to both standard transformer baselines and dedicated CoT models. With a parameter count of only 0.9B, our model achieves an 8-shot score of 48.2%, outperforming Llama 3.2 (1B) CoT baseline by 3.8 percentage points. Most notably, JEPA-Reasoner significantly surpasses larger models that rely on standard reasoning paradigms. Despite being approximately $4\times$ smaller than Gemma 3 (4B), our model exhibits a performance gain of nearly 10%. While the Recurrent Depth model (Huginn-0125) offers a strong baseline at 42.1%, it requires nearly four times the parameter count to achieve results that are still 6.1% lower than JEPA-Reasoner. Consequently, these results serve as strong empirical evidence of JEPA-Reasoner’s capability to handle complex natural language reasoning tasks, effectively applying mathematical logic to solve problems.
7 Theoretical Advantage of Decoupled Architecture
The robustness of our decoupled architecture stems from its ability to decouple the high-level reasoning process from low-level token generation. We can formalize this advantage by analyzing the probabilistic assumptions and information flow within the coupled and decoupled paradigms.
7.1 Model Dynamics and Probabilistic Factorization
Let $R=(r_{1},r_{2},\dots,r_{T})$ be the sequence of latent reasoning states and $X=(\hat{x}_{1},\hat{x}_{2},\dots,\hat{x}_{T})$ be the sequence of generated tokens.
Classical Transformer Model.
A standard Transformer model $\mathcal{M}_{t}$, implicitly defines a joint probability distribution that is factorized sequentially. The generation of the state and token at step $t$ depends on both the state and the sampled token from step t1t-1:
$P(R,X)=\prod_{t=1}^{T}P(r_{t},\hat{x}_{t}|r_{t-1},\hat{x}_{t-1})$
In this formulation, the distribution for the next reasoning state $r_{t}$ is directly conditioned on the previously sampled token $\hat{x}_{t-1}$. Consequently, a sampling error at step t1t-1 (i.e., $\hat{x}_{t-1}\neq x^{*}_{t-1}$) introduces a persistent error into the reasoning state trajectory. This error corrupts the foundation for all subsequent reasoning and generation steps, leading to compounding error.
COCONUT Model.
The COCONUT model $\mathcal{M}_{c}$ is a coupled model with latent generation ability. It generates latent reasoning tokens before producing final tokens (coconut). Limited by the coupled architecture, latent vectors and tokens are arranged in the same sequence $Z=(z_{1},z_{2},\dots,z_{N})$, where each element $z_{t}$ can be either a continuous latent vector $r_{t}$ or a discrete token $\hat{x}_{t}$. Despite the different output types, the model follows the same autoregressive principle: every new element is conditioned on all prior elements.
$P(Z)=\prod_{t=1}^{N}P(z_{t}|z_{<t})$
Let’s consider a generation length of $T_{1}$ for latent reasoning, followed by $T_{2}$ tokens. The process unfolds as follows:
  • For the latent reasoning steps ($t=1,\dots,T_{1}$), the model generates $z_{t}=r_{t}$, conditioning on the previous latent vectors $z_{<t}=(r_{1},\dots,r_{t-1})$.
  • For the token generation steps ($t=T_{1}+1,\dots,T_{1}+T_{2}$), the model generates $z_{t}=\hat{x}_{t-T_{1}}$, conditioning on the full history of all preceding latent vectors and tokens, $z_{<t}=(r_{1},\dots,r_{T_{1}},\hat{x}_{1},\dots,\hat{x}_{t-T_{1}-1})$.
The unified sequence is the model’s critical limitation. Suppose the model has finished generating its reasoning chain $(r_{1},\dots,r_{T_{1}})$ and the first error appeared at the $n^{th}$ token, $\hat{x}_{n}\neq x^{*}_{n}$. For the very next step, $t=T_{1}+n+1$, the model make prediction based on the history $(r_{1},\dots,r_{T_{1}},\hat{x}_{1},\dots,\hat{x}_{n})$. The erroneous token $\hat{x}_{n}$ is now an immutable part of the model’s context, corrupting every subsequent decision. The error propagation is direct and unavoidable because reasoning and generation are inextricably linked in the same autoregressive sequence.
Decoupled Model $\mathcal{M}_{d}$:
In contrast, our JEPA-Reasoner architecture imposes a structural constraint on the generative process, yielding a more robust factorization where the reasoning chain is generated independently of the token sampling:
$P(R,X)$ $=P(R)\cdot P(X|R)$
$=\left(\prod_{t=1}^{T}P(r_{t}|r_{t-1})\right)\cdot\left(\prod_{t=1}^{T}P(\hat{x}_{t}|R,\hat{x}_{1:t-1})\right)$
This factorization reveals two key theoretical advantages.
  1. 1.
    Error Containment: The reasoning trajectory’s probability, $P(R)$, is independent of the token generation process $P(X|R)$. An error in sampling a token $\hat{x}_{t-1}$ has no mathematical pathway to influence the reasoning trajectory $R$. The high-level plan remains intact and stable. Furthermore, the normalization of reasoning vectors $r_{t}$ to the unit hypersphere ensures this trajectory is inherently bounded, preventing error amplification within the reasoning dynamics itself.
  2. 2.
    Mechanism for Recovery: At every step $t$, the token generator $P(\hat{x}_{t}|\cdot)$ is conditioned on the entire, lossless reasoning chain $R$. This provides a strong, stable signal that allows the Talker to potentially recover from a local token error in its own history ($\hat{x}_{1:t-1}$), an effect empirically validated in our ablation study (Appendix C).
This inherent error containment and recovery mechanism explains the superior robustness observed in our CFG experiments (Section 5.2).
8 Summary
We introduce JEPA-Reasoner, a novel architecture that decouples latent space reasoning from token generation. Our approach enables continuous latent reasoning guidance while mitigating step-by-step error propagation. Efficient parallel training was also made possible without sacrificing latent reasoning performance compared with COCONUT. Our experiments on synthetic tasks suggest that by decoupling the high-level latent space reasoning process from low-level token generation, JEPA-Reasoner produced promising potential for multi-threaded reasoning and exhibited enhanced robustness to input noise and error accumulation when generating structured sequences.
Appendix A Data for Tree Search Experiment
The following is an example of data used in the tree-search experiment:
NL,NO,LJ,LI,OA,OM,JB,JH,IG,IF,AD,A
E,MK,MC[ROOT]N[TARGET]K[ROUTE]NOMK
Visualization of the example can be seen in Figure 3. In the sequence, each character pair represents a parent-child node pair, with the former one being the parent node and the later one being the child. All pairs are separated by a comma. The searching task is specified after the tree-structure definition, with special token [ROOT] indicating the tree root, [TARGET] indicating which leaf to search for, and [ROUTE] stating the correct searching route. All characters, comma, [ROOT], [TARGET], and [ROUTE] have a corresponding token, making up the whole vocabulary for the model along with the padding token and the end-of-sentence token.
Figure 3: Visualization of the tree structure in the given example.
Appendix B Further Details for CFG Experiments
B.1 CFG Rules and Sample
CFG can hierarchically produce highly structured expressions by replacing non-terminal symbols at each level with next-level symbols following a production rule, as shown in Figure 4. A sequence of terminal symbols is considered to be valid if it can be transformed back to the root symbol with dynamic programming and the given production rule. The recursive structure and local ambiguity of CFG sequences enable them to model the rich and recursive structure in languages, including grammar and logic. We designed our own CFG following the method used by (allen-physics). The production rule used in our experiments is a five-level CFG production rule set featuring three terminal symbols with 3 or 4 rule lengths at each level, which typically generates long (typically 600 to 700 symbols per sample) and locally ambiguous sequences. A visualization of the production rule used in this experiment can be seen in Figure 5.
Since even a 5-level CFG production rule that allows each non-terminal symbol to produce 2 to 3 symbols in the next level (simpler than our 5-level production rule that allows each non-terminal symbol to produce 3 to 4 symbols in the next level) is capable of producing more than $4\times 10^{8}$ distinctive sequences, we conclude that the models in the CFG experiments does not rely on memorizing possible sequences during training to achieve high accuracy on completion tasks.
Previous research (allen-physics) shows that Transformer blocks can encode the structure of CFG rules within parameters. We assume that a robust model should be able to recognize the high-level structure of the input sequence, thus ignoring faulty tokens in the input. Since each high-level element in our CFG sequence produces 3 to 4 tokens, the model should be able to maintain relatively stable performance across at least 4 generation steps.
Figure 4: A picture demonstrating how CFG sequence is generated. It involves replacing non-terminal symbols at each level with symbols from the next level according to a given rule.
Figure 5: CFG production rule used to generate training and test samples in Section 5.2. This rule gives sequence lengths ranging from about 600 symbols to 700 symbols.
B.2 A Sample CFG Sequence
We demonstrate a sample CFG sequence from the training dataset: 15 16 16 14......14 16 14 (666 terminal symbols in total). It consists of three kinds of terminal symbols.
B.3 Detailed Results of Token Level Error Test
We provide the accuracy of Transformer baseline and JEPA-Reasoner under different noise levels on CFG completion task mentioned in Section 5.2 here in Table 7. Note that the accuracy is the total correct symbols generated divided by the total symbols needed in the generation task.
Table 7: Absolute and relative (in parentheses, normalized to zero noise) accuracy of Transformer baseline and JEPA-Reasoner under different noise levels on CFG completion task
Noise Model Gen=1 Step=2 Step=3 Step=4
0.00 $T$ 0.9230 (1.00) 0.9181 (1.00) 0.6185 (1.00) 0.6861 (1.00)
$R$ 0.9191 (1.00) 0.7405 (1.00) 0.5722 (1.00) 0.6646 (1.00)
0.05 $T$ 0.8748 (0.95) 0.7867 (0.86) 0.5281 (0.85) 0.6140 (0.89)
$R$ 0.8506 (0.93) 0.6703 (0.91) 0.5242 (0.92) 0.6130 (0.92)
0.10 $T$ 0.7996 (0.87) 0.6958 (0.76) 0.4857 (0.79) 0.5600 (0.82)
$R$ 0.7688 (0.84) 0.6109 (0.82) 0.4797 (0.84) 0.5708 (0.86)
0.15 $T$ 0.7543 (0.82) 0.6273 (0.68) 0.4514 (0.73) 0.5273 (0.77)
$R$ 0.7256 (0.79) 0.5678 (0.77) 0.4581 (0.80) 0.5451 (0.82)
0.20 $T$ 0.7092 (0.77) 0.5868 (0.64) 0.4316 (0.70) 0.5005 (0.73)
$R$ 0.6797 (0.74) 0.5336 (0.72) 0.4428 (0.77) 0.5195 (0.78)
0.25 $T$ 0.6742 (0.73) 0.5643 (0.61) 0.4129 (0.67) 0.4704 (0.69)
$R$ 0.6426 (0.70) 0.5021 (0.68) 0.4152 (0.73) 0.5039 (0.76)
0.30 $T$ 0.6402 (0.69) 0.5442 (0.59) 0.4010 (0.65) 0.4521 (0.66)
$R$ 0.6061 (0.66) 0.4919 (0.66) 0.4201 (0.73) 0.4865 (0.73)
Appendix C Ablation Study of Talker Model
We conduct an ablation study to verify two critical properties of our decoupled architecture:
  1. 1.
    Reasoning Dominance: The reasoning process is strictly driven by the JEPA-Reasoner’s latent trajectory, preventing the Talker from bypassing the architecture to perform independent inference.
  2. 2.
    Linguistic Capability: Talker module acts as an effective “Language Interface,” capable of translating abstract latent vectors into grammatically correct and semantically coherent natural language.
We test this by corrupting the output of Reasoner in different ways. Our empirical experiments show strong evidence that the Talker cannot reason on its own and serves primarily as a readout mechanism for the Reasoner’s planning.
C.1 Experiment Setup
Using the training method mentioned in Section 4, we use the JEPA-Reasoner model initialized with Transformer blocks trained on C4 and Wikitext (wikitext-dataset) dataset in this experiment setup to produce a human-readable result. We conducted controlled experiments using two sample inputs from the training dataset to evaluate the dependency of the Talker model on the Reasoner’s output:
The first sample is: “Francis Bacon was an English philosopher and statesman who served as Attorney General and Lord Chancellor of England under King James I. Bacon argued for the importance of natural philosophy, guided by the scientific, his works remained influential”. This sample is used as JEPA-Reasoner and Talker’s input unless mentioned otherwise. The second sample is “Jean-Paul Sartre was a French philosopher, political activist, biographer, and literary critic. Sartre was one of the key figures in the philosophy of existentialism (and phenomenology).”, which is used in the “Semantic Mismatch” experiment
We systematically corrupted different components of the input to isolate the contribution of each part:
  • Baseline: Normal operation with clean Reasoner output
  • Random String Replacement: Replace Reasoner output with a random string
  • Initial Token Corruption: Keep Reasoner output clean, but replace Talker’s initial input string (decoder input) with a random string.
  • Gaussian Noise: Replace Reasoner output with Gaussian noise ($\mu=0$, $\sigma=1$)
  • Semantic Mismatch: Use Reasoner output from a different sentence
C.2 Results
Table 8 presents the results of our ablation experiments. The Talker’s initial input tokens were the first 10 tokens of the Francis Bacon sample (Francis Bacon was an English philosopher and statesman who in natural language across most experiments unless mentioned otherwise.
Table 8: Ablation Study Results: Talker Model Output Under Different Input Corruptions.
Condition Talker Output
Baseline Bacon was an English philosopher and statesman who served as Attorney General and Lord Chancellor of England under King James I. Bacon argued for the importance of natural philosophy, guided by the scientific; his works remained influential
Random String Francis Bacon was an English philosopher and statesman whoTGvIUujapaUYDIUbIyuTviuiYtvU
Random Initial Tokens SsDXicoundfdfiyx served as Attorney General and Lord Chancellor of England under King James I. Richardson argued for the importance of natural philosophy, guided by the scientific, his works remained influential
Gaussian Noise Francis Bacon was an English philosopher and statesman who 0 not deathsardumn technicalipt’tputalase2006 (ightsringst 0q virtually lowesteral 1 new date 2007 0 results;
Semantic Mismatch Francis Bacon was an English philosopher and statesman who, political activist, biographer, and literary critic. Zartre was one of the key figures in the philosophy of existentialism (and phenomenology).
Latent Dependency (Proof of No Independent Reasoning).
When the Reasoner output is replaced with Gaussian noise, the Talker produces incoherent output. This demonstrates that the Talker cannot generate meaningful content without proper latent representations.
Robustness to Token Errors (Proof of Linguistic Capability).
When only the initial input tokens are corrupted, the Talker can still produce largely coherent content, guided by the clean Reasoner output, although some localized errors occur (e.g., Richardson appears). This suggests the Reasoner’s latent representations carry the primary semantic information.
Semantic Fidelity (The Decoupling Verification).
When using the Reasoner output from the Jean-Paul Sartre sample as Dual-Talker’s input latent. Although the first 10 tokens from the Francis Bacon sample are used as the initial input tokens of decoder blocks in Dual-Talker, it rapidly shifts to generating the Jean-Paul Sartre content. This provides strong evidence that the Talker genuinely utilizes the semantic content encoded by the JEPA-Reasoner. Meanwhile, the fact that the generated output remains grammatically fluid despite this conflict proves that the Talker successfully handles the linguistic realization of the Reasoner’s abstract concepts.
Appendix D K Value in Scaled Cosine Distance Loss
We tested $k$ from 1 to 6. All these $k$ values could produce a basic SST outcome that exhibits reasoning behaviors stated in all previous sections. With a careful tuning of $k$, we observed a stable improvement in the tree-search problem as shown in Figure 6:
Figure 6: Changes of correct major contributor rate with training steps. Zoom in to see details.
We choose the correct major contributor (the correct next-step latent vector plays the most important role in current-step mixed latent vectors) rate as the metric since it directly relates to the correctness of future predictions. Considering that when $k=4$, the model gains the highest correct major contributor rate (zoom in to distinguish the line of $k=4$ from the line of $k=5$), we choose to continue our experiment with $k=4$.
Appendix E Visualization of Mixed Latent Vectors
Figure 7: PCA analysis of latent representations of different tree leaves in the tree search experiment.
To visualize that JEPA-Reasoner can produce mixed latent vectors, we gathered the embedding vectors and model predictions from our tree-search experiments in this section. We extracted embedding vectors from distinct tree leaves alongside the model’s output latent representations after one forward pass. Principal Component Analysis (PCA) was applied to the collected embeddings and model predictions, and the visualization focuses on the first two principal components (PC1 and PC2).
As demonstrated in the Figure 7, predicted latent vectors (blue points in the figure) form a continuous cloud within the space spanned by discrete vocabulary embeddings (red diamond shapes). This distribution supports the experiment results that they are the linear combinations of vocabulary embeddings. Also, the predicted vectors do not converge to singular vocabulary points, providing empirical evidence for the hypothesis that JEPA-Reasoner is capable of maintaining information from multiple possible choices rather than committing to a single answer.

### v2

@@
 While Joint-Embedding Predictive Architecture (JEPA) has emerged as a powerful architecture for learning rich latent representations, it fundamentally lacks generative abilities.
-Meanwhile, latent space reasoning attempts for Transformer models like COCONUT do improve performance, but they ultimately rely on token-by-token generation, which still accumulates compounding error and relies on context information to gain reasoning insights.
-To address these limitations, we propose JEPA-Reasoner, a novel JEPA model enhanced with generative ability that reasons in latent space.
-We augment it with a separate action-taker model, Talker, to produce human-readable sentences.
-Our approach demonstrates that decoupling latent space reasoning and token generation enables JEPA-Reasoner to produce mixed latent vectors that might lay the foundation for multi-threaded reasoning, while performing autoregressive generation with superior robustness to compounding error.
+Meanwhile, current latent reasoning models remain limited by the token-by-token generation paradigm, which suffers from compounding errors and heavy context dependency.
+To address these limitations, we proposed JEPA-Reasoner, a novel JEPA-based architecture enhanced with generative ability for latent reasoning.
+We augment this architecture with a separate action-talker model, Talker, to reconstruct human-readable text from latent representations produced by the JEPA-Reasoner.
+Our work demonstrated that decoupling latent-space reasoning from token production enables JEPA-Reasoner to produce mixed latent vectors, laying a foundation for multi-threaded reasoning and achieving superior robustness against compounding errors in autoregressive generation.
 Foundation Models, Latent Reasoning, ICML
 1 Introduction
-The Joint-Embedding Predictive Architecture (JEPA) (i-jepa) has demonstrated strong performance in learning semantic world representations, exhibiting superior world understanding ability compared with traditional end-to-end generative models that work in pixel or token space.
-By predicting abstract representations in latent space, the JEPA architecture is able to filter irrelevant details and preserve essential information needed for prediction (i-jepa).
+The Joint-Embedding Predictive Architecture (JEPA) (i-jepa) has demonstrated strong performance in learning semantic world representations, exhibiting a superior ability to understand the world compared with traditional end-to-end generative models that operate in pixel- or token-space.
+By predicting abstract representations in latent space, the JEPA architecture can filter irrelevant details and preserve essential information needed for prediction (i-jepa).
 Such architecture has been proven to be a viable approach for representation learning and foundation model development.
@@
 However, JEPA models are inherently non-generative (m3-jepa) because of their objective: filling missing information in the current state (i-jepa), rather than generating new content.
-Besides, the predictor often requires a detailed, predetermined target state or task instruction (e.g., V-JEPA 2-AC (v-jepa2)), which is often unavailable in tasks demanding long-term planning and step-by-step reasoning.
+Furthermore, the predictor often requires a detailed, predetermined target state or task instruction (e.g., V-JEPA 2-AC (v-jepa2)), which is typically unavailable in tasks that demand long-term planning and step-by-step reasoning.
 This feature limits the application of the broad knowledge in JEPA models to generative tasks.
@@
 Even if the predicted probabilities of tokens are weighted and combined, LLMs still cannot go beyond single-threaded reasoners (llms-reason-singel-thread).
-Although several previous research have explored latent space reasoning for Transformer models, the end-to-end generation goal of a single coupled model limited the full potential of latent space reasoning, while also making training complex and inefficient (coconut).
+Although several previous studies have explored latent space reasoning for Transformer models, the end-to-end generation goal of a single coupled model has limited the full potential of latent space reasoning and made training complex and inefficient (coconut).
 To address these limitations, we propose JEPA-Reasoner, a novel decoupled architecture that utilizes separate models for reasoning and token generation.
 The reasoning model, JEPA-Reasoner, transforms the JEPA framework from a target-conditioned system into an autoregressive generative model.
-Operating entirely within the continuous, normalized latent space, JEPA-Reasoner focuses solely on latent space reasoning, offloading the token generation task to its action-taker module: Talker.
+Operating entirely within the continuous, normalized latent space, JEPA-Reasoner focuses solely on latent-space reasoning, offloading the token-generation task to its action-taker module, Talker.
 In contrast to existing latent space reasoning solutions like COCONUT, the division of objective frees the reasoner from expression burden and enables continuous latent guidance (compared with non-continuous latent guidance that relies on context information to retrieve latent reasoning results (coconut)) during the token generation process.
-Our key insight is that performing pure reasoning in an abstract, continuous, and normalized latent space without token generation burden allows for the construction of high-level reasoning chains that carry rich semantic information throughout the autoregressive process while correctly ignoring irrelevant details or distracting information.
-Our empirical experiments show further improved abilities beyond that: by operating in the latent space, JEPA-Reasoner can maintain multiple hypotheses during the reasoning process simultaneously and mitigate the catastrophic error propagation associated with discrete token sampling.
+Our key insight is that performing pure reasoning in an abstract, continuous, and normalized latent space without the burden of token generation enables the construction of high-level reasoning chains that carry rich semantic information throughout the autoregressive process while correctly ignoring irrelevant details or distracting information.
+Our empirical experiments demonstrate further improvements beyond that: by operating in the latent space, JEPA-Reasoner can maintain multiple hypotheses simultaneously during the reasoning process and mitigate catastrophic error propagation associated with discrete token sampling.
 2 Related Work
@@
 However, these models are non-generative.
-Attempts to make JEPA generative, such as D-JEPA (d-jepa), utilize the learned representations to condition diffusion models for data generation (e.g., text to images, text to audio) but still failed to enable sequential reasoning or planning within the JEPA framework itself.
+Attempts to make JEPA generative, such as D-JEPA (d-jepa), use learned representations to condition diffusion models for data generation (e.g., text-to-image, text-to-audio), but they still fail to enable sequential reasoning or planning within the JEPA framework itself.
 In contrast, our key innovation is to adapt the core JEPA objective for autoregressive latent-space generation.
@@
 Besides, whether iterating over tokens or deepening the computation per token, both paradigms remain bound by causal masking, meaning early token generation cannot be guided by future reasoning states.
-Considering these limitations, our key innovation is to decouple latent space reasoning and token generation.
-By generating a complete latent chain first, we enable consistent latent guidance where all tokens are generated from the full reasoning trajectory.
-This produces an answer with higher quality that is less prone to error propagation.
+Given these limitations, our key innovation is to decouple latent-space reasoning and token generation.
+By generating a complete latent chain first, we enable consistent latent guidance, with all tokens generated from the whole reasoning trajectory.
+This produces a higher-quality answer that is less prone to error propagation.
 Additionally, the decoupled design enables efficient optimization with a single forward pass in latent space, unlike coupled models like COCONUT or Recurrent Depth transformers, which require multiple synchronized passes or complex recurrent unrolling.
@@
 Techniques like Chain-of-Thought (cot) improve reasoning by generating intermediate steps, but still operate at the token level.
-JEPA-Reasoner aims to improve robustness of autoregressive generation by moving the reasoning process into a continuous, abstract latent space, reducing the impact of localized errors.
+JEPA-Reasoner aims to enhance robustness of autoregressive generation by moving the reasoning process into a continuous, abstract latent space, reducing the impact of localized errors.
 3 Model Architecture
@@
 We utilize L2 normalization to prevent exploding magnitude caused by residual connections.
-In the modified Transformer block, we apply a non-learnable QK-Norm (qk-norm) to make it more numerically stable.
+In the modified Transformer block, we apply a non-learnable QK-Norm (qk-norm) to ensure numerical stability.
 Latent Space Generation.
@@
 Detailed information about the components of the two Talkers is shown in Table 1.
-Mono-Talker does not have an embedding layer or encoders, it only has decoders.
+Mono-Talker consists of decoders only, without an embedding layer or encoders.
 Mono-Talker is designed for reconstruction tasks that do not require context information, receiving latent vectors from JEPA-Reasoner and constructs the complete token sequence in one forward pass.
@@
 The training process of JEPA-Reasoner consists of two main phases.
-The first stage is pretraining that teaches basic knowledge (e.g., grammar and commonsense) to the model.
-The second stage is self-supervised training (SST), which adapts the model to perform consistent latent space reasoning.
+(1) Pretraining, to acquire linguistic competence and world knowledge, and (2) Self-Supervised Training (SST), to adapt the model for consistent reasoning in the continuous latent space.
 4.1 Pretraining
@@
 Objective and Methodology.
-The model is trained as a standard decoder-only Transformer on the next-token prediction task in a teacher-forcing way.
-We employ tied word embeddings that share the weight of the embedding layer with a temporary LM head.
-The LM head is only used in pretraining and is removed after pretraining is finished.
-The L2 normalization layer is disabled in the pretraining phase to enable simple and direct reuse of current Transformer training recipes.
-Considering that tied word embedding encourages $\mathbf{W}_{Embed}\cdot\mathbf{W}_{Embed}^{T}=I$ and $\mathbf{v}_{pred}\cdot\mathbf{v}_{embed}=\|\mathbf{v}_{pred}\|\cdot\|\mathbf{v}_{embed}\|\cdot\cos(\theta)$, the tied-embedding approach will indirectly encourage angular alignment between predicted vectors and embedding vectors, which facilitates the subsequent transition from token-level to latent-level prediction.
+The model is trained as a standard decoder-only Transformer on the next-token prediction task via teacher forcing.
+To facilitate the transition to the subsequent SST phase, we introduce specific architectural changes:
+  • Tied Embeddings: We share weights between the embedding layer and a temporary Language Model (LM) head. This LM head is discarded after pretraining.
+  • L2 Normalization: We disable the L2 normalization layer during this phase to maintain compatibility with standard Transformer training recipes.
+The use of tied embeddings also indirectly facilitates angular alignment between the predicted vectors and the embedding vectors since it encourages $\mathbf{W}_{Embed}\cdot\mathbf{W}_{Embed}^{T}=I$ and $\mathbf{v}_{pred}\cdot\mathbf{v}_{embed}=\|\mathbf{v}_{pred}\|\cdot\|\mathbf{v}_{embed}\|\cdot\cos(\theta)$.
+This implicit angular alignment creates a smoother transition from token-level prediction to latent-level prediction in the subsequent SST phase.
 4.2 Self-Supervised Training
-The SST phase adapts the pretrained model for making consistent predictions in the continuous latent space.
-Since the model is fully transforming into a latent space reasoner instead of a token generator, the ability to produce correct logits no longer matters.
-Considering this, we apply similar self-supervised training as Meta’s JEPA series (i-jepa) in this stage.
-Without the need for autoregressively generating final token outputs to compute a loss, self-supervised training enables efficient parallel training compared with COCONUT (coconut).
-Objective and Methodology.
-The temporary LM head in the pretraining stage was discarded and L2 normalization layer was restored.
-The model is now optimized to predict the latent representation of the next sequence segment, with a consistent dimensional semantics as what the embedding layer produces.
-We switch to scaled cosine distance loss, aligning with the L2 normalization strategy used to ensure stability during autoregressive looping and focusing the learning on angular differences:
-$\mathcal{L}(\theta,\theta^{\prime})=k-k\cdot\text{cos}(h_{\text{pred}}(\theta),h_{\text{target}}(\theta^{\prime}))$ (1)
-where $k$ is the scalar, $h_{\text{pred}}$ is the predicted latent vector by the Reasoner (parameters $\theta$), and $h_{\text{target}}$ is the target latent vector from the EMA encoder (parameters $\theta^{\prime}$).
-In our empirical tests, we find that normal cosine distance loss failed to support enough optimization when the loss is small.
-We tested a series of $k$ values, and chose $k=4$ in our experiments (Refer to Appendix D for more detail).
-Target Generation.
-The weight updating method of the target embedding layer is different from that of the input embedding layer.
-The input embedding layer always applies the latest weights, while the target embedding layer utilizes exponential moving average to generate its weights from the historical weights of the input embedding layer.
+The SST phase transforms the model from a token generator into a pure latent-space reasoner.
+We adopt a self-supervised approach inspired by the JEPA series (i-jepa).
+Unlike methods such as COCONUT (coconut), our approach avoids the computational overhead of autoregressively generating final tokens or looping latent vectors multiple times for loss computation, enabling more efficient parallel training.
+Optimization Objective.
+The model is now optimized to predict the latent representation of the next sequence segment.
+We discard the temporary LM head and restore the L2 normalization layer.
+We apply a similar self-supervised training strategy proposed in Meta’s JEPA series (i-jepa) and switch to scaled cosine distance loss, aligning with the L2 normalization strategy to ensure stability during autoregressive looping and encouraging the model to learn angular differences:
+$\mathcal{L}(\theta,\theta^{\prime})=k\cdot\left(1-\cos(h_{\text{pred}}(\theta),h_{\text{target}}(\theta^{\prime}))\right)$ (1)
+where $k$ is a scalar, $h_{\text{pred}}$ is the latent vector predicted by the Reasoner (parameters $\theta$) and $h_{\text{target}}$ is the target vector from the EMA encoder (parameters $\theta^{\prime}$).
+We adopt a scaling factor of $k=4$ as we empirically observed that standard cosine distance yields insufficient gradients when the loss is small.
+(See Appendix D for more discussions about scaled cosine distance loss).
+Target Generation (EMA).
+To prevent rank collapse while allowing angular adjustment, we employ an asymmetric weight update strategy:
+  1. 1.
+    Input Embedding (Online): Updated via standard backpropagation.
+  2. 2.
+    Target Embedding (Target): Updated via exponential moving average (EMA) of the input embedding weights.
 We applied a high momentum value of 0.98 to prevent rank collapse in the embedding layer while ensuring enough space to adjust for angular alignment.
@@
 5.2 Robustness to Error Perturbation
-In the following sections, we demonstrate that decoupling reasoning chain generation from token production enables superior robustness under noisy conditions.
-While coupled models must simultaneously maintain reasoning coherence and produce correct tokens, our decoupled approach allows the reasoning model to focus solely on maintaining logical consistency in latent space, exhibiting better generation quality.
-This section contains two experiments that focus on two different sources of errors: token level error and latent space noise.
+We demonstrate that decoupling reasoning generation from token production yields superior robustness.
+While coupled models must simultaneously maintain reasoning coherence and produce correct tokens, our decoupled approach allows the reasoning model to focus solely on logical consistency in the latent space, demonstrating superior generation quality and robustness to noisy input.
+We validate this advantage through two experiments targeting token-level errors and latent space noise, respectively.
 5.2.1 Experiment Methods
-Robustness Test for Token Level Error
+Robustness Test for Token Level Error.
 To evaluate the robustness of the decoupled model on token-level errors in the input sequence, we randomly replace 0% to 30% ground truth tokens in the input sequence with incorrect tokens.
@@
 using the exact match metric.
-Robustness Test for Latent Space Error
+Robustness Test for Latent Space Error.
 To evaluate the robustness of the JEPA-Reasoner model on perturbations in latent space, we compare the performance of JEPA-Reasoner and the coupled continuous reasoning model COCONUT on multi-step completion tasks.
-In this experiment, we let the COCONUT model autoregressively generate 4 latent vectors first, followed by 4 tokens.
-For JEPA-Reasoner, we simply let it generate 8 latent vectors and use the Talker module to reconstruct 8 tokens.
-For both models, we add Gaussian noise to the generated latent vector at each step with $\mu=0$ and $\sigma$ ranging from 0% to 15% of the maximum value in the model’s output.
-Accuracy is calculated across the last 4 tokens with the exact match metric.
+In this experiment, COCONUT generates four latent vectors followed by four tokens, whereas JEPA-Reasoner generates eight latent vectors and reconstructs eight tokens via the Talker module.
+We inject Gaussian noise at each step ($\mu=0$; $\sigma\in[0\%,15\%]$ relative to the maximum output value) and measure exact match accuracy on the final four tokens.
 5.2.2 Data Preparation
@@
 Total Blocks 24 24 24
-We first pretrain the Transformer models on CFG data using cross-entropy loss in token space.
-Since the pretraining methods of Transformer, COCONUT, and JEPA-Reasoner are identical, subsequent trainings are based on the same pretraining checkpoint [^2] Due to architectural differences, we initialize JEPA-Reasoner models using only the first $N$ blocks from the pretrained Transformer, where $N$ matches the JEPA-Reasoner’s block count..
-We conduct post-training to obtain the Transformer baseline $T$.
-COCONUT models are trained to first predict 4 hidden states, then generate 4 tokens.
-Cross-entropy loss is computed between the output logits and the target sequence, excluding the hidden state positions.
-We follow the training method mentioned in Section 4 to train the JEPA-Reasoner and Dual Talker model $R$.
-For all training processes mentioned above, we apply identical hyperparameters (with learning rate of $1\times 10^{-4}$, effective batch size of 128, and context length of 1024) and train until their loss stabilizes, then checkpoints of best performance were chosen as the representative.
+We first pretrain Transformer models on CFG data with cross-entropy loss, then pick the checkpoint with optimal accuracy on CFG completion task.
+This unified checkpoint serves as the initialization for all subsequent models.
+[^2] The JEPA-Reasoner is initialized using the first $N$ blocks of the pretrained Transformer to match its layer count.
+We then post-train the Transformer to obtain baseline $T$.
+COCONUT models are trained to predict a sequence of four hidden states followed by four tokens, with loss computed exclusively on the target tokens.
+The JEPA-Reasoner and Dual Talker model $R$ are trained following the methodology in Section 4.
+All training phases use identical hyperparameters (learning rate $1\times 10^{-4}$, effective batch size 128, context length 1024) and all models are trained until their loss stabilizes.
+The best-performing checkpoint is selected as the representative for each architecture.
 5.2.4 Results and Conclusions
@@
 The production rule used in our experiments is a five-level CFG production rule set featuring three terminal symbols with 3 or 4 rule lengths at each level, which typically generates long (typically 600 to 700 symbols per sample) and locally ambiguous sequences.
-A visualization of the production rule used in this experiment can be seen in Figure 5.
 Since even a 5-level CFG production rule that allows each non-terminal symbol to produce 2 to 3 symbols in the next level (simpler than our 5-level production rule that allows each non-terminal symbol to produce 3 to 4 symbols in the next level) is capable of producing more than $4\times 10^{8}$ distinctive sequences, we conclude that the models in the CFG experiments does not rely on memorizing possible sequences during training to achieve high accuracy on completion tasks.
@@
 Figure 4: A picture demonstrating how CFG sequence is generated. It involves replacing non-terminal symbols at each level with symbols from the next level according to a given rule.
-Figure 5: CFG production rule used to generate training and test samples in Section 5.2. This rule gives sequence lengths ranging from about 600 symbols to 700 symbols.
 B.2 A Sample CFG Sequence

### v3

@@
 Abstract
-While Joint-Embedding Predictive Architecture (JEPA) has emerged as a powerful architecture for learning rich latent representations, it fundamentally lacks generative abilities.
-Meanwhile, current latent reasoning models remain limited by the token-by-token generation paradigm, which suffers from compounding errors and heavy context dependency.
-To address these limitations, we proposed JEPA-Reasoner, a novel JEPA-based architecture enhanced with generative ability for latent reasoning.
-We augment this architecture with a separate action-talker model, Talker, to reconstruct human-readable text from latent representations produced by the JEPA-Reasoner.
-Our work demonstrated that decoupling latent-space reasoning from token production enables JEPA-Reasoner to produce mixed latent vectors, laying a foundation for multi-threaded reasoning and achieving superior robustness against compounding errors in autoregressive generation.
+Current autoregressive language models couple high-level reasoning and low-level token generation into a single sequential process, making the reasoning trajectory vulnerable to compounding expression errors.
+We propose JEPA-Reasoner, a novel architectural paradigm that decouples these tasks using a Joint-Embedding Predictive Architecture (JEPA) for pure latent-space reasoning and a separate Talker module for linguistic reconstruction.
+By isolating the reasoning engine from the discrete token-sampling process, our architecture enables: (1) Error Containment, where token-level failures cannot propagate into the latent reasoning chain; (2) Continuous Guidance, providing the generator with access to the entire lossless reasoning trajectory; and (3) Representation of Uncertainty, allowing the model to maintain multiple hypotheses via mixed latent vectors.
+Controlled experiments on synthetic and natural language tasks demonstrate that this decoupling enables a 0.9B model to achieve a 149.5% improvement in 8-shot GSM8K accuracy over a coupled Transformer baseline trained on identical data.
+This work shifts the focus from scaling coupled models to investigating decoupled architectures as a more robust foundation for complex reasoning.
 Foundation Models, Latent Reasoning, ICML
 1 Introduction
-The Joint-Embedding Predictive Architecture (JEPA) (i-jepa) has demonstrated strong performance in learning semantic world representations, exhibiting a superior ability to understand the world compared with traditional end-to-end generative models that operate in pixel- or token-space.
-By predicting abstract representations in latent space, the JEPA architecture can filter irrelevant details and preserve essential information needed for prediction (i-jepa).
-Such architecture has been proven to be a viable approach for representation learning and foundation model development.
-Various JEPA implementations, including I-JEPA (i-jepa), V-JEPA2 (v-jepa2), and M3-JEPA (m3-jepa), have shown success across various modalities and downstream tasks.
-However, JEPA models are inherently non-generative (m3-jepa) because of their objective: filling missing information in the current state (i-jepa), rather than generating new content.
-Furthermore, the predictor often requires a detailed, predetermined target state or task instruction (e.g., V-JEPA 2-AC (v-jepa2)), which is typically unavailable in tasks that demand long-term planning and step-by-step reasoning.
-This feature limits the application of the broad knowledge in JEPA models to generative tasks.
-Furthermore, while traditional token-level autoregressive models have sequential reasoning capabilities, their token-by-token generation process is prone to compounding errors.
-Even if the predicted probabilities of tokens are weighted and combined, LLMs still cannot go beyond single-threaded reasoners (llms-reason-singel-thread).
-Although several previous studies have explored latent space reasoning for Transformer models, the end-to-end generation goal of a single coupled model has limited the full potential of latent space reasoning and made training complex and inefficient (coconut).
-To address these limitations, we propose JEPA-Reasoner, a novel decoupled architecture that utilizes separate models for reasoning and token generation.
-The reasoning model, JEPA-Reasoner, transforms the JEPA framework from a target-conditioned system into an autoregressive generative model.
-Operating entirely within the continuous, normalized latent space, JEPA-Reasoner focuses solely on latent-space reasoning, offloading the token-generation task to its action-taker module, Talker.
-In contrast to existing latent space reasoning solutions like COCONUT, the division of objective frees the reasoner from expression burden and enables continuous latent guidance (compared with non-continuous latent guidance that relies on context information to retrieve latent reasoning results (coconut)) during the token generation process.
-Our key insight is that performing pure reasoning in an abstract, continuous, and normalized latent space without the burden of token generation enables the construction of high-level reasoning chains that carry rich semantic information throughout the autoregressive process while correctly ignoring irrelevant details or distracting information.
-Our empirical experiments demonstrate further improvements beyond that: by operating in the latent space, JEPA-Reasoner can maintain multiple hypotheses simultaneously during the reasoning process and mitigate catastrophic error propagation associated with discrete token sampling.
+The dominant paradigm for Large Language Models (LLMs) is a unified autoregressive process where internal reasoning and external expression are inextricably linked.
+In this coupled framework, the model must predict the next token based on a history that includes its own previously sampled discrete tokens.
+This creates a fundamental bottleneck: any localized error in token selection—whether a grammatical slip or a sub-optimal word choice—pollutes the context window, directly corrupting the trajectory of all subsequent reasoning.
+While recent efforts such as COCONUT (Hao et al., 2024) and Recurrent Depth models (Geiping et al., 2025) have explored latent-space reasoning, they remain largely constrained by this coupling.
+They either interleave latent vectors with tokens in a single sequence or rely on causal masking that prevents early tokens from being informed by future reasoning states.
+Furthermore, these models struggle to represent uncertainty, as the discrete nature of token-level coupling often forces a collapse into a single reasoning path (Wu et al., 2025).
+In this paper, we propose a novel decoupled reasoning paradigm: JEPA-Reasoner, an architecture that fully decouples latent-space reasoning from token generation.
+Our approach utilizes the Joint-Embedding Predictive Architecture (JEPA) (Assran et al., 2023) not merely for representation learning, but as an independent, autoregressive reasoning engine.
+The Reasoner operates entirely within a continuous, normalized latent space, generating a complete reasoning chain before any tokens are produced.
+A separate, lightweight Talker module then reconstructs human-readable text under the guidance of this completed latent trajectory.
+By isolating these functions, we demonstrate three emergent properties of the decoupled architecture:
+  1. 1.
+    Mathematical Error Containment: We provide a probabilistic factorization showing that because the reasoning chain is generated independently, token-sampling errors have no mathematical pathway to influence the logic of the Reasoner.
+  2. 2.
+    Architectural Efficiency: Our experiments show that JEPA-Reasoner utilizes internally learned knowledge more effectively than coupled baselines.
+When trained on identical datasets, the decoupled architecture overcomes the ”reasoning plateau” observed in small-scale models, showing superior in-context scaling from 5-shot to 8-shot tasks.
+  3. 3.
+    Mixed Latent States: We empirically validate that the Reasoner can produce ”mixed latent vectors” that encode multiple possible reasoning paths simultaneously, laying the foundation for multi-threaded reasoning without the overhead of beam search or multiple decodes.
+Rather than focusing on competitive SOTA benchmarks through massive scaling, this work serves as an architectural investigation into the benefits of decoupling.
+We show that even without sophisticated reinforcement learning or industrial training recipes, transitioning from a coupled to a decoupled paradigm yields significant gains in robustness and reasoning depth, suggesting a new path forward for foundation model design.
 2 Related Work
 Joint-Embedding Predictive Architectures (JEPA).
-JEPA (i-jepa) introduced a framework that makes predictions in representation space.
+JEPA (Assran et al., 2023) introduced a framework that makes predictions in representation space.
 It utilizes self-supervised training to learn latent states that are not directly human-readable.
 A predictor module was trained to predict the target state based on encoded inputs.
-Multiple variants of this architecture have extended the JEPA family to various modalities and downstream tasks (i-jepa).
+Multiple variants of this architecture have extended the JEPA family to various modalities and downstream tasks (Assran et al., 2023).
 However, these models are non-generative.
-Attempts to make JEPA generative, such as D-JEPA (d-jepa), use learned representations to condition diffusion models for data generation (e.g., text-to-image, text-to-audio), but they still fail to enable sequential reasoning or planning within the JEPA framework itself.
+Attempts to make JEPA generative, such as D-JEPA (Chen et al., 2025), use learned representations to condition diffusion models for data generation (e.g., text-to-image, text-to-audio), still fail to enable sequential reasoning or planning within the JEPA framework itself.
 In contrast, our key innovation is to adapt the core JEPA objective for autoregressive latent-space generation.
 Latent Space Reasoning.
-Previous work on latent space reasoning mainly focuses on looping hidden states, either through horizontal autoregression like COCONUT (coconut), or vertical recurrent depth scaling (recurrent-latent-reasoning).
+Previous work on latent space reasoning mainly focuses on looping hidden states, either through horizontal autoregression like COCONUT (Hao et al., 2024), or vertical recurrent depth scaling (Geiping et al., 2025).
 However, these paradigms utilize a single coupled model for both latent reasoning and token generation, ignoring the mismatch between the two tasks: one requires high-level global planning, decision-making, and choice tracing, while the other requires local correctness in grammar and fluency.
@@
 This produces a higher-quality answer that is less prone to error propagation.
-Additionally, the decoupled design enables efficient optimization with a single forward pass in latent space, unlike coupled models like COCONUT or Recurrent Depth transformers, which require multiple synchronized passes or complex recurrent unrolling.
+Additionally, the decoupled design enables efficient optimization with a single forward pass in latent space, unlike coupled models such as COCONUT or Recurrent Depth Transformers, which require multiple synchronized passes or complex recurrent unrolling.
 Autoregressive Models and Robustness.
@@
 While proven powerful on various tasks, this approach is known to suffer from compounding errors in long-horizon tasks.
-Techniques like Chain-of-Thought (cot) improve reasoning by generating intermediate steps, but still operate at the token level.
+Techniques like Chain-of-Thought (Wei et al., 2022) improve reasoning by generating intermediate steps, but still operate at the token level.
 JEPA-Reasoner aims to enhance robustness of autoregressive generation by moving the reasoning process into a continuous, abstract latent space, reducing the impact of localized errors.
@@
   • JEPA-Reasoner: Generate sequential latent space reasoning chains independently.
-  • Talker: Translates the latent states into tokens. Note that the Talker is not able to make predictions. Its task is to reconstruct tokens based entirely on the latent output from JEPA-Reasoner.
+  • Talker: Translates the latent states into tokens.
+Note that the Talker can not make predictions.
+(Refer to ablation study in Appendix C) Its task is to reconstruct tokens solely based on the latent reasoning chain produced by JEPA-Reasoner.
 3.1 JEPA-Reasoner
-Figure 1: Architecture of JEPA-Reasoner and its action taker (Talker).
+Figure 1: Architecture of JEPA-Reasoner and Talker.
 The Reasoner consists of an embedding layer as the token encoder and Transformer blocks as the predictor.
@@
 We utilize L2 normalization to prevent exploding magnitude caused by residual connections.
-In the modified Transformer block, we apply a non-learnable QK-Norm (qk-norm) to ensure numerical stability.
+In the modified Transformer block, we apply a non-learnable QK-Norm (Dehghani et al., 2023) to ensure numerical stability.
 Latent Space Generation.
-Unlike traditional JEPA models, in which the predictor is aimed at filling missing information in the current state (i-jepa), the predictor of JEPA-Reasoner generates the next latent matrix representing the subsequent reasoning steps.
+Unlike traditional JEPA models, in which the predictor is aimed at filling missing information in the current state (Assran et al., 2023), the predictor of JEPA-Reasoner generates the next latent matrix representing the subsequent reasoning steps.
 Crucially, this generated latent matrix is not projected into vocabulary probabilities via an LM head.
@@
 Following standard JEPA methodology, the target encoder weights are an exponential moving average (EMA) of the data encoder weights, providing stable and rich training targets.
-Given the normalized nature of our latent space, we use scaled cosine similarity loss computed entirely in latent space, ensuring the predictor learns consistent feature representations and dynamics (refer to Section 4.2 for more detail).
-3.2 Action-Taker Model
-In this experiment, there is only one action-taker model: Talker.
+Given the normalized nature of our latent space, we use scaled cosine similarity loss computed entirely in the latent space, ensuring the predictor learns consistent feature representations and dynamics (refer to Section 4.2 for more details).
+3.2 Talker Model
 The Talker model is a standard Transformer-based model trained independently.
@@
 The SST phase transforms the model from a token generator into a pure latent-space reasoner.
-We adopt a self-supervised approach inspired by the JEPA series (i-jepa).
-Unlike methods such as COCONUT (coconut), our approach avoids the computational overhead of autoregressively generating final tokens or looping latent vectors multiple times for loss computation, enabling more efficient parallel training.
+We adopt a self-supervised approach inspired by the JEPA series (Assran et al., 2023).
+Unlike methods such as COCONUT (Hao et al., 2024), our approach avoids the computational overhead of autoregressively generating final tokens or looping latent vectors multiple times for loss computation, enabling more efficient parallel training.
 Optimization Objective.
@@
 We discard the temporary LM head and restore the L2 normalization layer.
-We apply a similar self-supervised training strategy proposed in Meta’s JEPA series (i-jepa) and switch to scaled cosine distance loss, aligning with the L2 normalization strategy to ensure stability during autoregressive looping and encouraging the model to learn angular differences:
+We apply a similar self-supervised training strategy proposed in Meta’s JEPA series (Assran et al., 2023) and switch to scaled cosine distance loss, aligning with the L2 normalization strategy to ensure stability during autoregressive looping and encouraging the model to learn angular differences:
 $\mathcal{L}(\theta,\theta^{\prime})=k\cdot\left(1-\cos(h_{\text{pred}}(\theta),h_{\text{target}}(\theta^{\prime}))\right)$ (1)
@@
 We adopt a scaling factor of $k=4$ as we empirically observed that standard cosine distance yields insufficient gradients when the loss is small.
-(See Appendix D for more discussions about scaled cosine distance loss).
+(See Appendix D for detailed discussions about scaled cosine distance loss).
 Target Generation (EMA).
@@
 Based on this result, we examine the generated latent vectors to probe the reasoning behavior of JEPA-Reasoner.
+Visualization of Mixed Latent Vectors
+To visualize the mixed latent vectors, we extracted embedding vectors from distinct tree leaves alongside the model’s output latent representations after one forward pass.
+Principal Component Analysis (PCA) was applied to the collected embeddings and model predictions, and Figure 2 demonstrates the visualization of first two principal components (PC1 and PC2).
+Noticing that predicted latent vectors (blue points in the figure) form a continuous cloud within the space spanned by discrete vocabulary embeddings (red diamond shapes).
+This distribution supports the experiment results that they are the linear combinations of vocabulary embeddings.
+Also, the predicted vectors do not converge to singular vocabulary points, providing empirical evidence for the hypothesis that JEPA-Reasoner is capable of maintaining information from multiple possible choices rather than committing to a single answer.
+Figure 2: PCA analysis of latent representations of different tree leaves in the tree search experiment.
+Quantitative Analysis
 We calculated the distance from the predicted latent vector to the plane spanned by any two vocabulary vectors and sorted them from closest to farthest.
@@
 This discovery demonstrated that JEPA-Reasoner could make correct choices without completely discarding the other information that contains potentially correct choices.
-According to the previous COCONUT study (coconut), this behavior might lay the foundation for breadth-first multi-threaded reasoning.
+According to the previous COCONUT study (Hao et al., 2024), this behavior might lay the foundation for breadth-first multi-threaded reasoning.
 5.2 Robustness to Error Perturbation
 We demonstrate that decoupling reasoning generation from token production yields superior robustness.
-While coupled models must simultaneously maintain reasoning coherence and produce correct tokens, our decoupled approach allows the reasoning model to focus solely on logical consistency in the latent space, demonstrating superior generation quality and robustness to noisy input.
+While coupled models must simultaneously maintain reasoning coherence and produce correct tokens, our decoupled approach allows the reasoning model to focus solely on consistency in the latent space, demonstrating superior generation quality and robustness to noisy input.
 We validate this advantage through two experiments targeting token-level errors and latent space noise, respectively.
@@
 To evaluate the robustness of the decoupled model on token-level errors in the input sequence, we randomly replace 0% to 30% ground truth tokens in the input sequence with incorrect tokens.
-We compare the performance of JEPA-Reasoner and traditional Transformer models on multi-step completion tasks [^1] All scores obtained in these two tests are by testing the model across 5248 samples randomly chosen from the test dataset containing 100000 samples to minimize the bias introduced by randomness.
-using the exact match metric.
+We compare the JEPA-Reasoner against traditional Transformer models on multi-step completion tasks using the exact match metric.
+To minimize bias, we test on 5,248 samples randomly selected from a test set of 100,000.
 Robustness Test for Latent Space Error.
-To evaluate the robustness of the JEPA-Reasoner model on perturbations in latent space, we compare the performance of JEPA-Reasoner and the coupled continuous reasoning model COCONUT on multi-step completion tasks.
-In this experiment, COCONUT generates four latent vectors followed by four tokens, whereas JEPA-Reasoner generates eight latent vectors and reconstructs eight tokens via the Talker module.
-We inject Gaussian noise at each step ($\mu=0$; $\sigma\in[0\%,15\%]$ relative to the maximum output value) and measure exact match accuracy on the final four tokens.
+To evaluate JEPA-Reasoner’s robustness against perturbations in latent space, we compare its performance with the coupled continuous reasoning model COCONUT on multi-step completion tasks.
+In this setup, COCONUT generates four latent vectors followed by four tokens, whereas JEPA-Reasoner generates eight latent vectors and reconstructs eight tokens via the Talker module.
+We inject Gaussian noise at each step ($\mu=0$; $\sigma\in[0\%,15\%]$ relative to the maximum value in each model’s output latent vectors) and measure exact match accuracy on the final four tokens.
 5.2.2 Data Preparation
-Considering that the compounding error caused by different faulty tokens differs significantly, it is difficult to quantitatively analyze the model’s behavior under token-level errors (e.g., replacing keywords in the sentence will decrease the quality more considerably than replacing a word that functions as a connector).
-We followed previous work (allen-physics) and utilized Context-Free Grammar (CFG) for both experiments to create a controllable experiment setting.
-Our custom CFG production rule features three terminal symbols with rule lengths of 3 or 4.
-With this rule, we generated long (approximately 600 to 700 symbols) and complex sequences that require non-trivial work to solve.
-The complexity of the grammar ensures that high accuracy relies on learning the underlying structure of the CFG sequence rather than memorizing specific sequences (refer to Appendix B.1 for full CFG specifications and production methods).
+Quantifying and testing token-level robustness in natural language is challenging because the impact of error tokens varies significantly by token type (e.g., replacing a keyword is more damaging than replacing a connector).
+To create a controlled experiment setup, we followed previous work (Allen-Zhu and Li, 2023) and utilized Context-Free Grammar (CFG) for both experiments.
+Our custom CFG features three terminal symbols with rule lengths of 3 or 4, producing complex and long sequences of 600–700 symbols.
+The diversity of sequences generated ensures that high accuracy requires learning the underlying structure rather than simple memorization (see Appendix B for detailed discussion).
 5.2.3 Model Configurations and Training Methods
-We denote the vanilla Transformer baseline as $T$, the COCONUT-style coupled latent space reasoning model as $C$, and the decoupled model as $R$ (both JEPA-Reasoner and Talker are included).
+We define three model variants: the vanilla Transformer baseline ($T$), the coupled COCONUT model ($C$), and our decoupled JEPA-Reasoner ($R$).
 Table 3 shows detailed model configurations used in this section:
@@
 Total Blocks 24 24 24
-We first pretrain Transformer models on CFG data with cross-entropy loss, then pick the checkpoint with optimal accuracy on CFG completion task.
-This unified checkpoint serves as the initialization for all subsequent models.
-[^2] The JEPA-Reasoner is initialized using the first $N$ blocks of the pretrained Transformer to match its layer count.
-We then post-train the Transformer to obtain baseline $T$.
-COCONUT models are trained to predict a sequence of four hidden states followed by four tokens, with loss computed exclusively on the target tokens.
-The JEPA-Reasoner and Dual Talker model $R$ are trained following the methodology in Section 4.
-All training phases use identical hyperparameters (learning rate $1\times 10^{-4}$, effective batch size 128, context length 1024) and all models are trained until their loss stabilizes.
-The best-performing checkpoint is selected as the representative for each architecture.
+We first pretrain a Transformer on CFG data using cross-entropy loss.
+The checkpoint with the highest accuracy is selected as the initialization for all subsequent models  [^1] The JEPA-Reasoner is initialized using the first $N$ blocks of the pretrained Transformer to match its layer count..
+We then post-train $T$ as a baseline.
+$C$ is trained to predict a sequence of hidden states followed by tokens.
+$R$ is trained following the methodology in Section 4.
+All training phases use identical hyperparameters (learning rate $1\times 10^{-4}$, batch size 128, and context length 1024) and all models are trained until the loss stabilizes.
+Then the best-performing checkpoint is selected as the representative for each architecture.
 5.2.4 Results and Conclusions
-Our robustness evaluation demonstrates the advantages of the decoupled architecture.
-In the token-level error experiment, JEPA-Reasoner demonstrated less performance degradation when facing input noise during multi-step CFG completion tasks (Figure 2).
+The results of the above experiments confirm the robustness advantage of the decoupled architecture.
+In the token-level error experiment, JEPA-Reasoner demonstrated less performance degradation than the baseline ($T$) when facing input noise during multi-step CFG completion tasks (Figure 3).
 JEPA-Reasoner also exhibits higher performance across different magnitudes of Gaussian noise in the latent space error experiment (Table 4), providing more empirical evidence for its robustness advantage.
 Figure 3: The relative performance of JEPA-Reasoner $R$ and Transformer baseline $T$ at each generation step with different fractions of correct tokens in the input sequence being replaced with wrong ones.
-Table 4: Performance of JEPA-Reasoner $R$ and COCONUT $C$ under different latent noise levels.
+Table 4: Exact match accuracy of JEPA-Reasoner $R$ and COCONUT $C$ under different latent noise levels.
 $R$ $C$
@@
 $\sigma=0.15\times\max{(h_{t})}$ 0.4468 0.3629
-These results demonstrate that JEPA-Reasoner has the potential to address the limitations of existing paradigms in Section 2.
-By operating in a normalized latent space and offloading token generation to the Talker module, subsequent reasoning outputs do not condition on previous decisions, thereby mitigating error accumulation in the autoregressive process and enabling more robust sequential generation under noisy conditions.
+These results demonstrate that JEPA-Reasoner has the potential to address the limitations of existing paradigms discussed in Section 2.
+By operating in a normalized latent space and offloading token generation to the Talker module, subsequent reasoning outputs do not condition on previous decisions, thereby mitigating error accumulation in the autoregressive process and enabling robust sequential generation under noisy conditions.
 6 Real World Evaluation
-We trained a 694M JEPA-Reasoner and a 198M Mono-Talker on natural language using the training steps stated in Section 4. We analyze the performance in two ways:
+We evaluate the 694M JEPA-Reasoner and 198M Mono-Talker (0.9B combined) on natural language tasks following the training methods in Section 4. Our analysis focuses on two key comparisons:
   1. 1.
-    Performance difference between pretrained model (token level prediction) and JEPA-Reasoner (latent level prediction), paired with its Mono-Talker model.
+    Comparison with Token-Level Model: The performance difference between the fine-tuned Transformer model (token level prediction) and JEPA-Reasoner paired with its Mono-Talker model (latent level prediction).
   2. 2.
-    Performance difference between JEPA-Reasoner (paired with Mono-Talker) and other latent or non-latent reasoning models.
-6.1 Performance Gain After Latent Reasoning Adaptation
-We evaluated both the base Transformer model and JEPA-Reasoner on GSM8K (gsm8k). Table 5 shows the benchmark score under both 5-shot and 8-shot settings.
-Table 5: GSM8K benchmarks of base Transformer and JEPA-Reasoner
+    Comparison with Other Reasoning Paradigms: The performance of JEPA-Reasoner relative to other latent and non-latent reasoning models of similar scale.
+6.1 Model Training
+We first pretrain a standard 0.9B Transformer model on C4 and Wikitext (Merity et al., 2016) for 300k steps, then fine-tune it for 42k steps using general math QA pairs.
+The best-performing checkpoint is selected to initialize and further fine-tune a Transformer model to obtain the baseline for comparison.
+We also trained a JEPA-Reasoner + Talker pair following the methodology in Section 4 with the same data and hyperparameters used to obtain the Transformer baseline.
+The best-performing checkpoints of each architecture that emerged during each training phase are selected as their representative.
+6.2 Performance Gain from Decoupled Architecture
+We evaluated both the fine-tuned Transformer model and JEPA-Reasoner on the GSM8K (Cobbe et al., 2021) benchmark. Table 5 shows the performance under both 5-shot and 8-shot settings.
+Table 5: GSM8K performance of Transformer baseline and JEPA-Reasoner
 Model Accuracy (%)
 5-shot 8-shot
-Base Transformer 20.7 20.8
+Transformer Baseline 20.7 20.8
 JEPA-Reasoner + Talker 39.4 51.9
-Performance Gain.
-Comparing to the base Transformer, JEPA-Reasoner improves performance by 79.2% (5-shot) and 149.5% (8-shot), respectively, demonstrating that adapting to a decoupled latent space reasoning paradigm significantly increases the accuracy of expressing internally learned knowledge with explicitly stated token sequences.
-Considering that it takes 300k steps for pretraining but only 13k steps for SST, while using the same dataset, it shows great evidence that most of the performance gain came from the adaptation to latent reasoning, rather than SST steps.
-Also, the scaled cosine similarity loss used in SST encourages smooth latent representation transition rather than logical correctness, which further support that minimal new knowledge was learned during SST.
+Accuracy Gain.
+JEPA-Reasoner substantially outperforms the fine-tuned Transformer, achieving performance gains of 79.2% (5-shot) and 149.5% (8-shot) on the GSM8K benchmark compared with the baseline.
+These results confirm that adapting to a decoupled, latent-space reasoning paradigm significantly improves the model’s ability to utilize internally learned knowledge.
+Notably, both the Transformer baseline and JEPA-Reasoner are trained with identical data and hyperparameters to ensure fair comparison.
+Furthermore, the scaled cosine similarity loss used in SST prioritizes smooth latent transitions over logical correctness.
+Together, these factors suggest that the performance gain stems from the advantage of the decoupled latent reasoning architecture rather than new knowledge obtained during SST.
 In Context Learning.
-Crucially, while the Base Transformer’s performance stagnates between 5-shot and 8-shot settings (improving only by 0.1%), the JEPA-Reasoner utilizes the additional examples to nearly double its accuracy.
+Crucially, the JEPA-Reasoner exhibits superior scaling behavior.
+While the Base Transformer’s performance stagnates between 5-shot and 8-shot settings (improving only by 0.1%), the JEPA-Reasoner utilizes the additional examples to achieve a 131.7% improvement in performance.
 This scaling behavior indicates that the decoupled latent space reasoning architecture effectively overcomes the reasoning plateau, which was widely observed in small-scale token-based models, allowing for more robust logical deduction that is less constrained by surface-level token statistics.
-6.2 Performance Comparison with Other Reasoning Models
-We compare JEPA-Reasoner against standard Transformer models and other reasoning models.
-Table 6 shows the GSM8K benchmark performance of our JEPA-Reasoner model (paired with Talker) compared to other models of similar size or performance.
-Table 6: GSM8K benchmark comparison between JEPA-Reason and other reasoning models.
-Note that we could not find 8-shot GSM8K benchmark results for COCONUT models.
-Since COCONUT models sacrifice performance (compared to CoT) as a trade-off for efficiency (coconut), we did not include them in the performance comparison.
+6.3 Performance Comparison with Other Reasoning Paradigms
+We evaluate JEPA-Reasoner against established reasoning paradigms like Chain-of-Thought (CoT) and Recurrent Depth.
+Table 6 shows the GSM8K benchmark performance of our JEPA-Reasoner model (paired with Talker) relative to other models of comparable size or performance [^2] Scores for Llama 3.2-1B (Meta AI, 2024) and Huginn-0125 (Geiping et al., 2025) are cited from their official reports.
+As official 8-shot scores are unavailable for Qwen 3 0.6B, we evaluated it using lm-evaluation-harness (Gao et al., 2024), employing the same configuration used for JEPA-Reasoner..
+Comparison of JEPA-Reasoner against other reasoning paradigms on the GSM8K benchmark.
+Note that the COCONUT model is excluded as public 8-shot results are unavailable, and prior work indicates it trades performance for efficiency relative to CoT (Hao et al., 2024).
+Table 6: Comparison of JEPA-Reasoner against other reasoning paradigms on the GSM8K benchmark.
+Note that the COCONUT model is excluded as neither its checkpoints nor its 8-shot GSM8K benchmark result are publicly available, and prior work indicates it trades performance for efficiency relative to CoT (Hao et al., 2024).
 Reasoning Paradigm Example Model Size 8-shot Accuracy (%)
-Standard Gemma 3 4B 38.4
 CoT Llama 3.2 1B 44.4
@@
 Recurrent Depth Huginn-0125 3.5B 42.1
-Ours JEPA-Reasoner 0.9B 51.9
+Decoupled Latent Reasoning JEPA-Reasoner (Ours) 0.9B 51.9
 Analysis of Results.
-As illustrated in Table 6, JEPA-Reasoner demonstrates superior performance on the GSM8K benchmark compared to both standard transformer baselines and dedicated CoT models.
-With a parameter count of only 0.9B, our model achieves an 8-shot score of 48.2%, outperforming Llama 3.2 (1B) CoT baseline by 3.8 percentage points.
-Most notably, JEPA-Reasoner significantly surpasses larger models that rely on standard reasoning paradigms.
-Despite being approximately $4\times$ smaller than Gemma 3 (4B), our model exhibits a performance gain of nearly 10%.
-While the Recurrent Depth model (Huginn-0125) offers a strong baseline at 42.1%, it requires nearly four times the parameter count to achieve results that are still 6.1% lower than JEPA-Reasoner.
-Consequently, these results serve as strong empirical evidence of JEPA-Reasoner’s capability to handle complex natural language reasoning tasks, effectively applying mathematical logic to solve problems.
+As illustrated in Table 6, JEPA-Reasoner demonstrates superior performance on the GSM8K benchmark compared to both CoT models and other latent reasoning architectures.
+With a parameter count of only 0.9B, our model achieves an 8-shot score of 51.9%, outperforming the CoT baseline of the same size (Llama 3.2-1B) by 7.5 percentage points.
+While the Recurrent Depth model (Huginn-0125) offers a strong baseline at 42.1%, it requires nearly four times the parameter count to achieve results that are still 9.8 percentage points lower than JEPA-Reasoner.
+Consequently, these results serve as strong empirical evidence that JEPA-Reasoner’s decoupled latent reasoning architecture is capable of handling complex natural language reasoning tasks both effectively and efficiently.
 7 Theoretical Advantage of Decoupled Architecture
-The robustness of our decoupled architecture stems from its ability to decouple the high-level reasoning process from low-level token generation.
+the robustness advantage of JEPA-Reasoner is built on two complementary mechanisms: the decoupled architecture design and the normalized property of latent space.
 We can formalize this advantage by analyzing the probabilistic assumptions and information flow within the coupled and decoupled paradigms.
-7.1 Model Dynamics and Probabilistic Factorization
-Let $R=(r_{1},r_{2},\dots,r_{T})$ be the sequence of latent reasoning states and $X=(\hat{x}_{1},\hat{x}_{2},\dots,\hat{x}_{T})$ be the sequence of generated tokens.
-Classical Transformer Model.
-A standard Transformer model $\mathcal{M}_{t}$, implicitly defines a joint probability distribution that is factorized sequentially.
-The generation of the state and token at step $t$ depends on both the state and the sampled token from step t1t-1:
-$P(R,X)=\prod_{t=1}^{T}P(r_{t},\hat{x}_{t}|r_{t-1},\hat{x}_{t-1})$
-In this formulation, the distribution for the next reasoning state $r_{t}$ is directly conditioned on the previously sampled token $\hat{x}_{t-1}$.
-Consequently, a sampling error at step t1t-1 (i.e., $\hat{x}_{t-1}\neq x^{*}_{t-1}$) introduces a persistent error into the reasoning state trajectory.
-This error corrupts the foundation for all subsequent reasoning and generation steps, leading to compounding error.
-COCONUT Model.
-The COCONUT model $\mathcal{M}_{c}$ is a coupled model with latent generation ability.
-It generates latent reasoning tokens before producing final tokens (coconut).
-Limited by the coupled architecture, latent vectors and tokens are arranged in the same sequence $Z=(z_{1},z_{2},\dots,z_{N})$, where each element $z_{t}$ can be either a continuous latent vector $r_{t}$ or a discrete token $\hat{x}_{t}$.
-Despite the different output types, the model follows the same autoregressive principle: every new element is conditioned on all prior elements.
-$P(Z)=\prod_{t=1}^{N}P(z_{t}|z_{<t})$
-Let’s consider a generation length of $T_{1}$ for latent reasoning, followed by $T_{2}$ tokens. The process unfolds as follows:
-  • For the latent reasoning steps ($t=1,\dots,T_{1}$), the model generates $z_{t}=r_{t}$, conditioning on the previous latent vectors $z_{<t}=(r_{1},\dots,r_{t-1})$.
-  • For the token generation steps ($t=T_{1}+1,\dots,T_{1}+T_{2}$), the model generates $z_{t}=\hat{x}_{t-T_{1}}$, conditioning on the full history of all preceding latent vectors and tokens, $z_{<t}=(r_{1},\dots,r_{T_{1}},\hat{x}_{1},\dots,\hat{x}_{t-T_{1}-1})$.
-The unified sequence is the model’s critical limitation.
-Suppose the model has finished generating its reasoning chain $(r_{1},\dots,r_{T_{1}})$ and the first error appeared at the $n^{th}$ token, $\hat{x}_{n}\neq x^{*}_{n}$.
-For the very next step, $t=T_{1}+n+1$, the model make prediction based on the history $(r_{1},\dots,r_{T_{1}},\hat{x}_{1},\dots,\hat{x}_{n})$.
-The erroneous token $\hat{x}_{n}$ is now an immutable part of the model’s context, corrupting every subsequent decision.
-The error propagation is direct and unavoidable because reasoning and generation are inextricably linked in the same autoregressive sequence.
-Decoupled Model $\mathcal{M}_{d}$:
-In contrast, our JEPA-Reasoner architecture imposes a structural constraint on the generative process, yielding a more robust factorization where the reasoning chain is generated independently of the token sampling:
-$P(R,X)$ $=P(R)\cdot P(X|R)$
-$=\left(\prod_{t=1}^{T}P(r_{t}|r_{t-1})\right)\cdot\left(\prod_{t=1}^{T}P(\hat{x}_{t}|R,\hat{x}_{1:t-1})\right)$
-This factorization reveals two key theoretical advantages.
-  1. 1.
-    Error Containment: The reasoning trajectory’s probability, $P(R)$, is independent of the token generation process $P(X|R)$.
-An error in sampling a token $\hat{x}_{t-1}$ has no mathematical pathway to influence the reasoning trajectory $R$.
-The high-level plan remains intact and stable.
-Furthermore, the normalization of reasoning vectors $r_{t}$ to the unit hypersphere ensures this trajectory is inherently bounded, preventing error amplification within the reasoning dynamics itself.
-  2. 2.
-    Mechanism for Recovery: At every step $t$, the token generator $P(\hat{x}_{t}|\cdot)$ is conditioned on the entire, lossless reasoning chain $R$.
-This provides a strong, stable signal that allows the Talker to potentially recover from a local token error in its own history ($\hat{x}_{1:t-1}$), an effect empirically validated in our ablation study (Appendix C).
-This inherent error containment and recovery mechanism explains the superior robustness observed in our CFG experiments (Section 5.2).
+Denote $R=(r_{1},r_{2},\dots,r_{T})$ as the sequence of latent reasoning states and $X=(\hat{x}_{1},\hat{x}_{2},\dots,\hat{x}_{T})$ as the sequence of generated tokens.
+7.1 Decoupled Architecture Stops the Propagation of Error
+Assume that the joint probability can be factored into $P(R,X)=P(R)\cdot P(X|R)$. The autoregressive update rules for the reasoning state $r_{t}$ and token $\hat{x}_{t}$ at step $t$ are defined as:
+$\begin{cases}r_{t}=f_{\theta}(r_{t-1})&\text{Reasoner Update}\\ \hat{x}_{t}\sim g_{\phi}(r_{t},\hat{x}_{1:t-1})&\text{Talker Sampling}\end{cases}$
+where $f_{\theta}$ is the deterministic predictor function of the Reasoner and $g_{\phi}$ is the probabilistic decoding function of the Talker.
+We define the sensitivity of the future reasoning state $r_{t+1}$ with respect to the current generated token $\hat{x}_{t}$ using partial derivatives as follows.
+Noticing that by the definition of decoupled architecture, $r_{t}$ depends solely on $r_{t-1}$.
+Considering that the generated token $\hat{x}_{t}$ does not appear in the input argument of $f_{\theta}$.
+Which means that $\frac{\partial r_{t}}{\partial\hat{x}_{\tau}}=0\quad\forall\tau\geq 1$.
+Using the chain rule, we obtain:
+$\frac{\partial r_{t+1}}{\partial\hat{x}_{t}}=\frac{\partial f_{\theta}(r_{t})}{\partial r_{t}}\cdot\underbrace{\frac{\partial r_{t}}{\partial\hat{x}_{t}}}_{0}=\mathbf{0}$
+As a conclusion, there is no mathematical pathway for token-space errors produced during sampling to propagate into the latent reasoning trajectory, which enables better robustness during generation
+7.2 Bounded Error via Normalization and Angular Alignment
+In this section, we demonstrate that unlike standard autoregressive models where errors can compound in both magnitude and direction, JEPA-Reasoner’s error bound is explicitly enforced by the L2 normalization layer and implicitly minimized by the angular training objective.
+This dual constraint prevents the reasoning trajectory produced from diverging unbounded, providing the theoretical basis for the superior robustness observed in our experiments.
+Explicitly Constrained by Architecture
+By constraining all latent states to the unit hypersphere $\mathcal{S}^{d-1}$, the divergence between an arbitrary trajectory produced by JEPA-Reasoner during runtime (denoted by $R$) and an optimal reference trajectory (denoted by $R^{*}$) is strictly bounded.
+Since the predictor output is passed through an L2 normalization layer before looping back, every reasoning state $r_{t}$ satisfies $\|r_{t}\|_{2}=1$.
+Consequently, the Euclidean distance between the predicted state and the optimal state is physically constrained by the diameter of the hypersphere:
+$\|r_{t}-r_{t}^{\prime}\|_{2}\leq\text{diam}(S^{d-1})=2$
+Implicitly Minimized by Training Objective
+Furthermore, the scaled cosine distance loss $\mathcal{L}_{t}\propto 1-\cos(r_{t},r^{*}_{t})=\frac{1}{2}\|r_{t}-r^{*}_{t}\|_{2}^{2}$ used in our Self-Supervised Training (SST) phase encourages the model to be aligned with a continuous trajectory on a bounded hypersphere instead of making leaps between discrete tokens like a traditional Transformer, implicitly reducing the risk of producing arbitrary latent vectors that diverges from the optimal reasoning path.
 8 Summary
@@
 Efficient parallel training was also made possible without sacrificing latent reasoning performance compared with COCONUT.
-Our experiments on synthetic tasks suggest that by decoupling the high-level latent space reasoning process from low-level token generation, JEPA-Reasoner produced promising potential for multi-threaded reasoning and exhibited enhanced robustness to input noise and error accumulation when generating structured sequences.
+Our experiments suggest that by decoupling the high-level latent space reasoning process from low-level token generation, JEPA-Reasoner produced promising potential for multi-threaded reasoning and exhibited enhanced robustness to input noise and error accumulation when generating structured sequences.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning.
+There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
 Appendix A Data for Tree Search Experiment
@@
 The recursive structure and local ambiguity of CFG sequences enable them to model the rich and recursive structure in languages, including grammar and logic.
-We designed our own CFG following the method used by (allen-physics).
+We designed our own CFG following the method used by (Allen-Zhu and Li, 2023).
 The production rule used in our experiments is a five-level CFG production rule set featuring three terminal symbols with 3 or 4 rule lengths at each level, which typically generates long (typically 600 to 700 symbols per sample) and locally ambiguous sequences.
 Since even a 5-level CFG production rule that allows each non-terminal symbol to produce 2 to 3 symbols in the next level (simpler than our 5-level production rule that allows each non-terminal symbol to produce 3 to 4 symbols in the next level) is capable of producing more than $4\times 10^{8}$ distinctive sequences, we conclude that the models in the CFG experiments does not rely on memorizing possible sequences during training to achieve high accuracy on completion tasks.
-Previous research (allen-physics) shows that Transformer blocks can encode the structure of CFG rules within parameters.
+Previous research (Allen-Zhu and Li, 2023) shows that Transformer blocks can encode the structure of CFG rules within parameters.
 We assume that a robust model should be able to recognize the high-level structure of the input sequence, thus ignoring faulty tokens in the input.
@@
 C.1 Experiment Setup
-Using the training method mentioned in Section 4, we use the JEPA-Reasoner model initialized with Transformer blocks trained on C4 and Wikitext (wikitext-dataset) dataset in this experiment setup to produce a human-readable result.
+Using the training method mentioned in Section 4, we use the JEPA-Reasoner model initialized with Transformer blocks trained on C4 and Wikitext (Merity et al., 2016) dataset in this experiment setup to produce a human-readable result.
 We conducted controlled experiments using two sample inputs from the training dataset to evaluate the dependency of the Talker model on the Reasoner’s output:
@@
 Considering that when $k=4$, the model gains the highest correct major contributor rate (zoom in to distinguish the line of $k=4$ from the line of $k=5$), we choose to continue our experiment with $k=4$.
-Appendix E Visualization of Mixed Latent Vectors
-Figure 6: PCA analysis of latent representations of different tree leaves in the tree search experiment.
-To visualize that JEPA-Reasoner can produce mixed latent vectors, we gathered the embedding vectors and model predictions from our tree-search experiments in this section.
-We extracted embedding vectors from distinct tree leaves alongside the model’s output latent representations after one forward pass.
-Principal Component Analysis (PCA) was applied to the collected embeddings and model predictions, and the visualization focuses on the first two principal components (PC1 and PC2).
-As demonstrated in the Figure 6, predicted latent vectors (blue points in the figure) form a continuous cloud within the space spanned by discrete vocabulary embeddings (red diamond shapes).
-This distribution supports the experiment results that they are the linear combinations of vocabulary embeddings.
-Also, the predicted vectors do not converge to singular vocabulary points, providing empirical evidence for the hypothesis that JEPA-Reasoner is capable of maintaining information from multiple possible choices rather than committing to a single answer.